# Parameter Symmetry Breaking and Restoration Determines the Hierarchical Learning in AI Systems

### v1

@@ Abstract
The dynamics of learning in modern large AI systems is hierarchical, often characterized by abrupt, qualitative shifts akin to phase transitions observed in physical systems. While these phenomena hold promise for uncovering the mechanisms behind neural networks and language models, existing theories remain fragmented, addressing specific cases. In this paper, we posit that parameter symmetry breaking and restoration serve as a unifying mechanism underlying these behaviors. We synthesize prior observations and show how this mechanism explains three distinct hierarchies in neural networks: learning dynamics, model complexity, and representation formation. By connecting these hierarchies, we highlight symmetry – a cornerstone of theoretical physics – as a potential fundamental principle in modern AI.
@@ 1 Introduction
More and more phenomena that are virtually universal in the learning process have been discovered in contemporary AI systems. These phenomena are shared by models with different architectures, trained on different datasets, and with different training techniques. The existence of these universal phenomena calls for one or a few universal explanations. However, until today, most of the phenomena are instead described by narrow theories tailored to explain each phenomenon separately – often focusing on specific models trained on specific tasks or loss functions and in isolation from other interesting phenomena that are indispensable parts of the deep learning phenomenology. Certainly, it is desirable to have a universal perspective, if not a universal theory, that explains as many phenomena as possible. In the spirit of science, a universal perspective should be independent of system details such as variations in minor architecture definitions, choice of loss functions, training techniques, etc. A universal theory would give the field a simplified paradigm for thinking about and understanding AI systems and a potential design principle for a new generation of more efficient and capable models.
@@
  • Hierarchy of Learning Dynamics: distinct temporal regimes arise during learning – this includes abrupt complexity jumps (Simon et al., 2023; Jacot et al., 2021; Abbe et al., 2023), progressive sharpening and flattening (Cohen et al., 2021), and beyond-linear dynamics of training (Zhu et al., 2022);
@@
  • Hierarchy of Model Complexity: the functional complexity of models adapts to the target function – this is exhibited in simplicity biases (Kalimeris et al., 2019), compressive coding through the information bottleneck (Tishby et al., 2000; Tishby & Zaslavsky, 2015), and the “blessing of dimensionality” in overparameterized nets (Galanti et al., 2021; Zhang et al., 2017);
@@
  • Hierarchy of Neural Representation: distinct spatial structures arise in the layers of neural networks, with progressively deeper layers tending to encode increasingly abstract information – this is evidenced in the structured representations such as neural collapse (Papyan et al., 2020), hierarchical encoding of features (Zeiler & Fergus, 2014), and universal alignment of representations across models (Huh et al., 2024).
@@
Figure 1: The division of solution space into hierarchies given by distinct parameter symmetries. Left: Example solution space of a model with parameter symmetries can be divided into hierarchies with boundaries prescribed by symmetry-breaking conditions. The more symmetry there is, the more restricted the hypothesis space becomes. Middle: The temporal (learning dynamics) and spatial (layer-wise information processing) dynamics of AI models can be characterized by the transitions between different symmetry groups. The solid line shows a symmetry restoration dynamics when the parameter transitions from a low-symmetry state to a high one either through time or through layers. The dashed lines show compositional dynamics where the model follows two symmetry breaking and then a symmetry restoration. Actual learning dynamics of neural networks may involve the model first learning by breaking symmetries before regularization effects dominate (Li et al., 2021), restoring symmetry. Similarly, for spatial processing, neural networks are found to break symmetries in early layers and restore symmetries in final layers (Section 5). Right: The more symmetry a model has, the more spatial, temporal, and functional hierarchies it has. Changes in symmetries can induce transitions between these hierarchies.
In this paper, we posit that these three seemingly unrelated hierarchies of deep learning phenomena can be collectively understood as arising from a single concept regarding the behavior of model parameter symmetries during training:
Main Hypothesis:
@@
The hierarchies of learning dynamics, model complexity, and representation formation in AI systems are primarily determined by parameter symmetry breaking and restoration.
@@
Prior art has certainly hinted at this idea, but here we draw together many separate threads to complete a single picture, as illustrated in Figure 1. The main idea is that parameter symmetry (defined in Section 2) divides the solution space into hierarchies, such that three key effects result:
@@
Given parameter symmetry as a unifying mechanism for deep learning, our understanding of the principles of AI can advance with explicit identification and engineering of symmetries in models, loss functions, and data. With the recent advances in how arbitrary parameter symmetries may be deliberately introduced or removed (Section 6), it is now possible to design symmetries matching practitioner intentions. Moreover, given widespread beliefs that hidden hierarchy phenomena of the kinds described above are beneficial, it follows one should be able to design models with intentional symmetries to introduce hierarchies into learning dynamics, the learned function complexities, and the layered structure of neural networks.
The next section introduces parameter symmetry. We then discuss existing supports for the three specific hypotheses given above. Section 6 discusses mechanisms that cause and methods for controlling symmetry breaking and restoration. The last section discusses alternative views. Formal theorems, experimental details, and additional figures are presented in the appendix.
@@
permutation $f(u,w)=f(w,u)$ identical neurons fully connected layers, ensembles
double rotation $f(U,W)=f(UA,A^{-1}W)$ low-rank solutions self-attention, matrix factorization
@@ 2 Parameter Symmetry in Deep Learning
$P_{G}$ is the projection matrix to the symmetry-invariant subspace of $G$ (Feit, 1982). Note that if $P_{G}\theta=\theta$, then $g\theta=\theta$ for any $g$ because $gP_{G}=P_{G}$. A striking fact for those unfamiliar with parameter symmetries is that many such parameter symmetries exist in widely used loss functions and neural networks (see Table 1). Understanding the implications of such symmetries has been an active area of research (Simsek et al., 2021; Entezari et al., 2021; Dinh et al., 2017; Saxe et al., 2013; Neyshabur et al., 2014; Tibshirani, 2021; Ioffe & Szegedy, 2015; Zhao et al., 2023b; Godfrey et al., 2022), and this has been challenging since the number of groups induced by these symmetries often grows exponentially in the size of the model. Many prior works have explored the effect of parameter symmetry on constraining the model complexity or learning dynamics (Li et al., 2016; Saxe et al., 2013; Du et al., 2018; Tanaka & Kunin, 2021; Zhao et al., 2022; Marcotte et al., 2023; Li et al., 2020). Because parameter symmetry naturally means that part of the solution space is redundant, having parameter symmetry can be seen as a form of “overparameterization.”
@@
Consider the parameter symmetries in the self-attention logit of a transformer, which can be written as (using notation from Vaswani et al. (2017))
$a_{ij}(W_{Q},W_{K})=X_{i}^{T}W_{Q}W_{K}X_{j}.\vspace{-1mm}$ (1)
@@
where $W_{Q}^{l}$ is the $l$-th column of $W_{Q}$ and $W_{K}^{l}$ is the $l$-th row of $W_{K}$. This means that it also has the permutation symmetry such that for any $l$ and $l^{\prime}$, $a_{ij}(W_{Q}^{l},W_{K}^{l},W_{Q}^{l^{\prime}},W_{K}^{l^{\prime}})=a_{ij}(W_{Q}^{l^{\prime}},W_{K}^{l^{\prime}},W_{Q}^{l},W_{K}^{l})$. An alternative way to see this is that permutation symmetry is a subgroup of the double rotation symmetry (because permutation matrices are invertible). For each $l$, there is also the sign flip symmetry, $a_{ij}(W_{Q}^{l},W_{K}^{l})=a_{ij}(-W_{Q}^{l},-W_{K}^{l})$, and the rescaling symmetry, $a_{ij}(\lambda W_{Q}^{l},\lambda^{-1}W_{K}^{l})$. Note that if $X$ is low-rank in the $z$-direction, these matrices also have a translation symmetry, where $a_{ij}(W_{K}^{l}+\lambda z)=a_{ij}(W_{K}^{l})$ for any $l$ and $\lambda\in\mathbb{R}$. Therefore, we have seen that even for a single self-attention layer, there are many symmetries. Note that these symmetries are independent of the data distribution and apply to every pair of $X_{i}$ and $X_{j}$ – this is the main reason these symmetries could significantly affect training neural networks. Recent works have shown that double rotation symmetry causes the self-attention layers to have a low-rank bias (Kobayashi et al., 2024).
@@ 3 Learning Dynamics is Symmetry-to- Symmetry
Symmetry has a direct influence on the loss landscape, and it thus affects the learning dynamics of neural networks through its effect on the landscape (Tatro et al., 2020; Lim et al., 2024). One primary effect of symmetry on the loss landscape is that it creates extended saddle points from which SGD or GD cannot escape (Li et al., 2016; Chen et al., 2023). Meanwhile, it has been found that when a neural network is initialized with small norm weights, its learning dynamics is primarily saddle-to-saddle (Jacot et al., 2021). In fact, neural networks have been found to converge often to saddle points (Alain et al., 2019; Ziyin et al., 2023). Given that symmetries are the primary origins of saddle points, it is natural to hypothesize that the learning dynamics of neural networks are not only saddle-to-saddle but symmetry-to-symmetry:
@@
Figure 2: DNN learning dynamics is symmetry-to-symmetry. Recent works suggested the learning of neural networks is primarily saddle-to-saddle (Jacot et al., 2021), and escaping these saddle points coincides with a sudden change in the complexity of the network (Abbe et al., 2023). At the same time, symmetries have been found to be the primary causes of the saddle points (Li et al., 2016; Ziyin, 2024). Once the symmetry is removed, saddle points seem to have disappeared when interpolating different solutions of the model (Lim et al., 2024). The figure shows that the loss jumps when symmetry breaking happens (black dotted lines) and plateaus when there is no symmetry breaking.
@@
Figure 2 shows an example where we train an MLP in a teacher-student setting using SGD. We initialize the model at a small norm, so the model’s initial state is approximately in the symmetric state, where the average pairwise distance is of order $10^{-4}$. The black dashed lines show when a pairwise distance exceeds $\Delta^{G}_{\rm th}$, signaling a symmetry breaking. We see that the symmetry-breaking times coincide well with the periods where the network learns rapidly. This shows that the saddle-to-saddle learning dynamics is potentially caused by the symmetry-to-symmetry transitions during training. Examples with polynomial and sinusoidal nets are shown in Section A.2.
Decomposability into Quadratic Models
@@
Another interesting point worth raising about symmetries is that if the model is close to a symmetric state $\theta^{*}$, then it can be decomposed into a sum of a quadratic model and a strictly smaller model. Ziyin (2024) showed that for a small $\Delta^{G}$,
$f(\theta,x)=f(\theta^{*})+\theta_{G}^{T}\nabla^{2}f(\theta^{*},x)\theta_{G}+o(\Delta^{G}),\vspace{-1mm}$ (4)
@@
where $\theta^{*}=P_{G}\theta$ is essentially a parameter projected to a $d-{\rm rank}(P_{G})$ subspace, and $\theta_{G}=(I-P_{G})\theta$ are the parameters projected to the symmetry breaking direction. Therefore, the second term can be seen as a quadratic model trained on a data transformation kernel of the form $\nabla^{2}f(x)$. This means that close to symmetry states, the learning dynamics must be “diligent” and cannot be characterized by NTK (Jacot et al., 2018) or lazy training (Chizat et al., 2018). This implies that parameter symmetry also plays a significant role in feature learning (Yang & Hu, 2020). In contrast, if a model has an antisymmetry, its expansion will only have odd order terms, and one may expect that such a model primarily behaves like a linear model.
@@
Figure 3: The complexity and generalization error of neural networks do not grow with width. A well-known observation in deep learning is that overparameterized networks not only work well, but also their generalization errors are empirically found to be essentially independent of width (Li & Wang, 2018; Pinto et al., 2024; Galanti et al., 2023; Mingard et al., 2025), an observation at odds with conventional bounds based on the Rademacher complexity (Zhang et al., 2017). The existence of parameter symmetries may solve this problem because, with a fixed regularization, the maximum surviving neurons are upper bounded by a constant. The left figure shows that Imagenet-pretrained ViT-Base (80M) and ViT-Large (300M) have similar degrees of symmetry in their self-attention layers. Here, each dot is a self-attention layer. The middle figure shows that the weight rank does not grow with increasing model size. A mechanism for this is that permutation symmetry implies that neuron weights of distance $o(\gamma)$ to each other must collapse – this means that within a fixed $n$-sphere, there can be at most $1/\gamma^{n}$ different neurons. This filling procedure is illustrated in the right figure: The orange circle denotes the parameter space, and the little circles are the neuron weights.
@@ 4 Symmetry Adaptively Limits Model Complexity
Another important implication of symmetries is that they control the effective number of parameters. When a system is symmetry-broken, new degrees of freedom usually emerge (known as a Goldstone mode in physics (Peskin, 2018)), and the system becomes higher-dimensional and more complex. When it is symmetry-restored, the effective dimension is reduced. This means that the symmetry boundaries naturally correspond to boundaries of different hierarchies of model capacities. This idea has a place in machine learning. For example, equivariant networks can improve the sample efficiency of training, and existing equivariant networks almost always involve introducing new parameter symmetry to the neural network (Maron et al., 2018; Bronstein et al., 2021). In Bayesian learning, parameter symmetry (and its generalizations) have been found to directly determine the generalization scaling of the model (Watanabe & Opper, 2010). Prior works also indicate how neural networks may break symmetries to adapt to different tasks (Fok et al., 2017).
More broadly, combining the idea that symmetry classes have different model complexity and the common observation that SGD tends to learn a function whose complexity is proportional to the complexity of the target function (Kalimeris et al., 2019; Mingard et al., 2025), it is natural to arrive at the following hypothesis:
@@
(Informal, (Ziyin et al., 2025b)) If the loss function has $G$-symmetry, and the initial $\theta\in\mathbb{R}^{d}$ is $G$-symmetric, then (1) there exists a model with $d-{\rm rank}(P_{G})$ parameters whose learning dynamics is the same as $\theta$, and (2), in the lazy training regime, this is equivalent to applying a rank $d-{\rm rank}(P_{G})$ mask to the NTK features.
Thus, symmetric solutions are low-capacity states from which gradient-based training methods cannot escape. More importantly, these symmetric solutions are preferred solutions when weight decay is used (Ziyin, 2024) or if the minibatch noise is strong due to a mechanism called “stochastic collapse” (Chen et al., 2023).
@@
In other words, parameter symmetries in the model may function like an Occam’s razor, where simpler and low-complexity solutions are favored – an observation that has been made across almost all modern neural models, whose generalization performances are independent of the size of the model (Zhang et al., 2017; Galanti et al., 2023), which is in discrepancy with common generalization bounds that predict a $\sqrt{\psi}$ deterioration with respect to the width $\psi$ (Neyshabur et al., 2018). One can test this hypothesis simply by training an MLP and a small transformer on different randomly generated teacher networks – see Section A.5. For different tasks, the same network converges to solutions with different symmetry classes corresponding to different levels of complexity. Similarly, for a fixed task, networks with different levels of capacity converge to similar degrees of symmetry. See Figure 3-Left for vision transformers (ViT) (Dosovitskiy, 2020) trained on Imagenet.
Here, we raise an insightful conjecture about the quantification of complexity control due to permutation symmetry, where $\gamma$ is the weight decay, $\eta/S$ is the learning-rate-batch-size ratio. Recall that the distance between two neurons is $\Delta_{ij}$.
@@
(Space Quantization Conjecture, Informal) In every layer with permutation symmetry, for two nonidentical neurons $i$ and $j$, $\Delta_{ij}>O(\kappa^{\beta})$ after training, where $\beta>0$ and $\kappa$ is the regularization strength.
We prove a special case in Section B.3. A primary mechanism for this absolute upper limit is a combination of regularization and symmetry (Section 6): (1) using weight decay $\gamma$, the model parameters must lie within a ball of radius $O(1/\gamma)$ to the origin; (2) with symmetry any neuron whose weights are with a distance of $\gamma$ to each other must become identical. This means that the number of surviving neurons is upper bounded by the number of $n-$hyperspheres of radius $O(\gamma)$ that one can fit in a hypersphere of radius $O(1/\gamma)$; this number can be upper bounded by $O(\gamma^{-n})$. This exponential index can be improved to be independent of $n$ and so does not suffer the curse of dimensionality (Section B.3). From a physics picture, the neurons start to form a “lattice” in the parameter space with the lattice distance $\gamma$.
Thus, when regularized, there are, at most, a finite number of nonidentical neurons within a layer, however wide it is. This implies that the actual complexity of the trained network must decrease as a function polynomial in the regularization strength. See Figure 3-Middle, where we train a two-hidden-layer network with varying widths. When weight decay or stochastic training is used, we see that the dimension of the latent representation saturates at a large model size. This could imply that the actual model complexity is much smaller than its apparent dimension and may be a key mechanism for understanding why highly overparametrized networks can still generalize. For example,  Bartlett et al. (2019) showed that the VC dimension of a two-layer piecewise linear network is $O(\psi)$. Our result implies that the nonidentical number of neurons must be $O(1)$ after training. This, in turn, means that the VC dimension of the model is independent of width.

### v2

# Parameter Symmetry Potentially Unifies Deep Learning Theory

@@ Abstract
The dynamics of learning in modern large AI systems is hierarchical, often characterized by abrupt, qualitative shifts akin to phase transitions observed in physical systems. While these phenomena hold promise for uncovering the mechanisms behind neural networks and language models, existing theories remain fragmented, addressing specific cases. In this position paper, we advocate for the crucial role of the research direction of parameter symmetries in unifying these fragmented theories. This position is founded on a centralizing hypothesis for this direction: parameter symmetry breaking and restoration are the unifying mechanisms underlying the hierarchical learning behavior of AI models. We synthesize prior observations and theories to argue that this direction of research could lead to a unified understanding of three distinct hierarchies in neural networks: learning dynamics, model complexity, and representation formation. By connecting these hierarchies, our position paper elevates symmetry – a cornerstone of theoretical physics – to become a potential fundamental principle in modern AI.
@@ 1 Introduction
More and more universal emergent phenomena of learning have been discovered in contemporary AI systems, and there is an urgent need for a unified theory to understand and explain them. These phenomena are shared by models with different architectures, trained on different datasets, and with different training techniques. The existence of these universal phenomena calls for one or a few universal explanations. However, until today, most of the phenomena are instead described by narrow theories tailored to explain each phenomenon separately – often focusing on specific models trained on specific tasks or loss functions and in isolation from other interesting phenomena that are indispensable parts of the deep learning phenomenology. Certainly, it is desirable to have a universal perspective, if not a universal theory, that explains as many phenomena as possible. In the spirit of science, a universal perspective should be independent of system details such as variations in minor architecture definitions, choice of loss functions, training techniques, etc. A universal theory would give the field a simplified paradigm for thinking about and understanding AI systems and a potential design principle for a new generation of more efficient and capable models.
@@
  • Hierarchy of Learning Dynamics: distinct temporal regimes arise during learning – this includes abrupt complexity jumps [61, 33, 1], progressive sharpening and flattening [14], and beyond-linear dynamics of training [82];
@@
  • Hierarchy of Model Complexity: the functional complexity of models adapts to the target function – this is exhibited in simplicity biases [34], compressive coding through the information bottleneck [67, 68], and the “blessing of dimensionality” in overparameterized nets [25, 78];
@@
  • Hierarchy of Neural Representation: distinct spatial structures arise in the layers of neural networks, with progressively deeper layers tending to encode increasingly abstract information – this is evidenced in the structured representations such as neural collapse [54], hierarchical encoding of features [77], and universal alignment of representations across models [30].
@@
Figure 1: The division of solution space into hierarchies given by distinct parameter symmetries. Left: Example solution space of a model with parameter symmetries can be divided into hierarchies with boundaries prescribed by symmetry-breaking conditions. The more symmetry there is, the more restricted the hypothesis space becomes. Middle: The temporal (learning dynamics) and spatial (layer-wise information processing) dynamics of AI models can be characterized by the transitions between different symmetry groups. The solid line shows a symmetry restoration dynamics when the parameter transitions from a low-symmetry state to a high one (through time or through layers). The dashed lines show a compositional dynamics where the model follows two symmetry breaking and then a restoration. Actual learning dynamics of neural networks may involve the model first learning by breaking symmetries before regularization effects dominate [43], restoring symmetry. Similarly, for spatial processing, neural networks are found to break symmetries in early layers and restore symmetries in final layers (Section 5). Right: The more symmetry a model has, the more spatial, temporal, and functional hierarchies it has. Changes in symmetries can induce transitions between these hierarchies.
In this paper, we argue for the position that the parameter symmetries should be a crucial perspective and direction of research in order to build universal theories of AI because, as we will collect key prior results to demonstrate, these three seemingly unrelated hierarchies of deep learning phenomena can be collectively understood as arising from having parameter symmetries in the model or loss function. Our central position is:
Position:
@@
Unifying theories and phenomenology of AI systems requires parameter symmetries, because the hierarchies of learning dynamics, model complexity, and representation formation are primarily determined by parameter symmetry breaking and restoration.
@@
Prior art has certainly hinted at this idea, but here we draw together many separate threads to complete a single picture, as illustrated in Figure 1. The main position is supported by three specific hypotheses, each of which is now extensively supported by prior results:
@@
If parameter symmetry is really a unifying mechanism for deep learning, our understanding of the principles of AI can advance with explicit identification and engineering of symmetries in models, loss functions, and data. With the recent advances in how arbitrary parameter symmetries may be deliberately introduced or removed (Section 6), it is now possible to design symmetries matching practitioner intentions. Moreover, given widespread beliefs that hidden hierarchy phenomena of the kinds described above are beneficial, it follows one should be able to design models with intentional symmetries to introduce hierarchies into learning dynamics, the learned function complexities, and the layered structure of neural networks.
The next section introduces parameter symmetry. We then discuss existing supports for the three specific hypotheses given above. Section 6 discusses mechanisms that cause and methods for controlling symmetry breaking and restoration. The last section discusses alternative views. Proofs, experimental details, andadditional figures are presented in the appendix.
@@
permutation $f(u,w)=f(w,u)$ identical neurons dense layers, ensembles
double rotation $f(U,W)=f(UA,A^{-1}W)$ low-rank solutions self-attention, linear nets
@@ 2 Parameter Symmetry in Deep Learning
These definitions are adopted from Ref. [88]. $P_{G}$ is the projection matrix to the symmetry-invariant subspace of $G$ [21]. Note that if $P_{G}\theta=\theta$, then $g\theta=\theta$ for any $g$ because $gP_{G}=P_{G}$. A striking fact for those unfamiliar with parameter symmetries is that many such parameter symmetries exist in widely used loss functions and neural networks (see Table 1). Understanding the implications of such symmetries has been an active area of research [62, 20, 16, 60, 51, 66, 31, 81, 29], and this has been challenging since the number of groups induced by these symmetries often grows exponentially in the size of the model. Many prior works have explored the effect of parameter symmetry on constraining the model complexity or learning dynamics [41, 60, 18, 64, 79, 45, 42]. Because parameter symmetry naturally means that part of the solution space is redundant, having parameter symmetry can be seen as a form of “overparameterization.”
@@
Consider the parameter symmetries in the self-attention logit of a transformer, which can be written as (using notation from Ref. [69])
$a_{ij}(W_{Q},W_{K})=X_{i}^{T}W_{Q}W_{K}X_{j}.$ (1)
@@
where $W_{Q}^{l}$ is the $l$-th column of $W_{Q}$ and $W_{K}^{l}$ is the $l$-th row of $W_{K}$. This means that it also has the permutation symmetry such that for any $l$ and $l^{\prime}$, $a_{ij}(W_{Q}^{l},W_{K}^{l},W_{Q}^{l^{\prime}},W_{K}^{l^{\prime}})=a_{ij}(W_{Q}^{l^{\prime}},W_{K}^{l^{\prime}},W_{Q}^{l},W_{K}^{l})$. An alternative way to see this is that permutation symmetry is a subgroup of the double rotation symmetry (because permutation matrices are invertible). For each $l$, there is also the sign flip symmetry, $a_{ij}(W_{Q}^{l},W_{K}^{l})=a_{ij}(-W_{Q}^{l},-W_{K}^{l})$, and the rescaling symmetry, $a_{ij}(\lambda W_{Q}^{l},\lambda^{-1}W_{K}^{l})$. Note that if $X$ is low-rank in the $z$-direction, these matrices also have a translation symmetry, where $a_{ij}(W_{K}^{l}+\lambda z)=a_{ij}(W_{K}^{l})$ for any $l$ and $\lambda\in\mathbb{R}$. Therefore, we have seen that even for a single self-attention layer, there are many symmetries. Note that these symmetries are independent of the data distribution and apply to every pair of $X_{i}$ and $X_{j}$ – this is the main reason these symmetries could significantly affect training neural networks. Recent works have shown that double rotation symmetry causes the self-attention layers to have a low-rank bias [35].
@@ 3 Learning Dynamics is Symmetry-to-Symmetry
Symmetry has a direct influence on the loss landscape, and it thus affects the learning dynamics of neural networks through its effect on the landscape [65, 44]. One primary effect of symmetry on the loss landscape is that it creates extended saddle points from which SGD or GD cannot escape [41, 12]. Meanwhile, it has been found that when a neural network is initialized with small norm weights, its learning dynamics is primarily saddle-to-saddle [33]. In fact, neural networks have been found to converge often to saddle points [3, 85]. Given that symmetries are the primary origins of saddle points, it is natural to hypothesize that the learning dynamics of neural networks are not only saddle-to-saddle but symmetry-to-symmetry:
@@
Figure 2: DNN learning dynamics is symmetry-to-symmetry. Recent works suggested the learning of neural networks is primarily saddle-to-saddle [33], and escaping these saddle points coincides with a sudden change in the complexity of the network [1]. At the same time, symmetries have been found to be the primary causes of the saddle points [41, 83]. Once the symmetry is removed, saddle points seem to have disappeared when interpolating different solutions of the model [44]. The figure repeats an experiment that is similar to those in Ref. [1] and shows that the loss jumps when symmetry breaking happens (black dotted lines) and plateaus when there is no symmetry breaking for the smallest init. As the init. scale becomes larger; such plateaus disappear because they are far away from symmetric states.
@@
Figure 2 shows an example where we train an MLP in a teacher-student setting using SGD. The model exhibits multiple symmetry jumps when we initialize the model at a small norm, i.e. the model’s initial state is approximately in the symmetric state. The black dashed lines show when a pairwise distance exceeds $\Delta^{G}_{\rm th}$, signaling a symmetry breaking. We see that the symmetry-breaking times coincide well with the periods where the network learns rapidly. This shows that the saddle-to-saddle learning dynamics is potentially caused by the symmetry-to-symmetry transitions during training. Examples with polynomial and sinusoidal nets are shown in Section A.2.
Decomposability and Feature Learning
@@
Another interesting point worth raising about symmetries is that if the model is close to a symmetric state $\theta^{*}$, then it can be decomposed into a sum of a quadratic model and a strictly smaller model. Ref. [83] showed that for a small $\Delta^{G}$,
$f(\theta,x)=f(\theta^{*})+\theta_{G}^{T}\nabla^{2}f(\theta^{*},x)\theta_{G}+o(\Delta^{G}),$ (4)
@@
where $\theta^{*}=P_{G}\theta$ is essentially a parameter projected to a $d-{\rm rank}(P_{G})$ subspace, and $\theta_{G}=(I-P_{G})\theta$ are the parameters projected to the symmetry breaking direction. Therefore, the second term can be seen as a quadratic model trained on a data transformation kernel of the form $\nabla^{2}f(x)$. This means that close to symmetry states, the learning dynamics must be “diligent” and cannot be characterized by NTK [32] or lazy training [13]. Thus, parameter symmetry also plays a significant role in feature learning [76] and a lot of work on neural network training with small initialization reflect the roles of symmetry in feature learning [60, 37, 28, 7, 61, 74]. In contrast, if a model is antisymmetric, its expansion will only have odd-order terms, and one expects that this model primarily behaves like a linear model.
@@
Figure 3: The complexity and generalization error of neural networks do not grow with width. A well-known observation in deep learning is that overparameterized networks not only work well, but also their generalization errors are empirically found to be essentially independent of width [40, 56, 26, 49], an observation at odds with conventional bounds based on the Rademacher complexity [78]. The existence of parameter symmetries may solve this problem because, with a fixed regularization, the maximum surviving neurons are upper bounded by a constant. The left figure takes publically available Imagenet-pretrained ViT-Base (80M) and ViT-Large (300M) have similar degrees of symmetry in their self-attention layers. Here, each dot is a self-attention layer. The middle figure shows that the weight rank does not grow with increasing model size. A mechanism for this is that permutation symmetry implies that neuron weights of distance $o(\gamma)$ to each other must collapse [73] – this means that within a fixed $n$-sphere, there can be at most $1/\gamma^{n}$ different neurons. This filling procedure is illustrated in the right figure: The orange circle denotes the parameter space, and the little circles are the neuron weights.
@@ 4 Symmetry Adaptively Limits Model Complexity
Another important implication of symmetries is that they control the effective number of parameters. When a system is symmetry-broken, new degrees of freedom usually emerge (known as a Goldstone mode in physics [55]), and the system becomes higher-dimensional and more complex. When it is symmetry-restored, the effective dimension is reduced. This means that the symmetry boundaries naturally correspond to boundaries of different hierarchies of model capacities. This idea has a place in machine learning. For example, equivariant networks can improve the sample efficiency of training, and existing equivariant networks almost always involve introducing new parameter symmetry to the neural network [46, 11]. In Bayesian learning, parameter symmetry (and its generalizations) have been found to directly determine the generalization scaling of the model [70]. Prior works also indicate how neural networks may break symmetries to adapt to different tasks [22].
More broadly, combining the idea that symmetry classes have different model complexity and the common observation that SGD tends to learn a function whose complexity is proportional to the complexity of the target function [34, 49], it is natural to arrive at the following hypothesis:
@@
(Informal, [88]) If the loss function has $G$-symmetry, and the initial $\theta\in\mathbb{R}^{d}$ is $G$-symmetric, then (1) there exists a model with $d-{\rm rank}(P_{G})$ parameters whose learning dynamics is the same as $\theta$, and (2), in the lazy training regime, this is equivalent to applying a rank $d-{\rm rank}(P_{G})$ mask to the NTK features.
Thus, symmetric solutions are low-capacity states from which gradient-based training methods cannot escape. More importantly, these symmetric solutions are preferred solutions when weight decay is used [83] or if the minibatch noise is strong due to a mechanism called “stochastic collapse” [12].
@@
In other words, parameter symmetries in the model may function like an Occam’s razor, where simpler and low-complexity solutions are favored – an observation that has been made across almost all modern neural models, whose generalization performances are independent of the size of the model [78, 26], which is in discrepancy with common generalization bounds that predict a $\sqrt{\psi}$ deterioration with respect to the width $\psi$ [50]. One can test this hypothesis simply by training an MLP and a small transformer on different randomly generated teacher networks – see Section A.5. For different tasks, the same network converges to solutions with different symmetry classes corresponding to different levels of complexity. Similarly, for a fixed task, networks with different levels of capacity converge to similar degrees of symmetry. See Figure 3-Left for vision transformers (ViT) [17] trained on Imagenet.
Here, one can raise an insightful conjecture about the quantification of complexity control due to permutation symmetry, where $\gamma$ is the weight decay, $\eta/S$ is the learning-rate-batch-size ratio. Recall that the distance between two neurons is $\Delta_{ij}$.
@@
(Space Quantization Conjecture [73], Informal) In every layer with permutation symmetry, for two nonidentical neurons $i$ and $j$, $\Delta_{ij}>O(\kappa^{\beta})$ after training, where $\beta>0$ and $\kappa$ is the regularization strength.
For completeness, a special case is proved in Section B.2. A primary mechanism for this absolute upper limit is a combination of regularization and symmetry (Section 6): (1) using weight decay $\gamma$, the model parameters must lie within a ball of radius $O(1/\gamma)$ to the origin; (2) with symmetry any neuron whose weights are with a distance of $\gamma$ to each other must become identical. This means that the number of surviving neurons is upper bounded by the number of $n-$hyperspheres of radius $O(\gamma)$ that one can fit in a hypersphere of radius $O(1/\gamma)$; this number can be upper bounded by $O(\gamma^{-n})$. This exponential index can be improved to be independent of $n$ and so does not suffer the curse of dimensionality (Section B.2). From a physics picture, the neurons start to form a “lattice” in the parameter space with the lattice distance $\gamma$.
Thus, when regularized, there are, at most, a finite number of nonidentical neurons within a layer, however wide it is. This implies that the actual complexity of the trained network must decrease as a function polynomial in the regularization strength. See Figure 3-Middle, where we train a two-hidden-layer network with varying widths. When weight decay or stochastic training is used, we see that the dimension of the latent representation saturates at a large model size. This could imply that the actual model complexity is much smaller than its apparent dimension and may be a key mechanism for understanding why highly overparametrized networks can still generalize. For example, Ref.  [9] showed that the VC dimension of a two-layer piecewise linear network is $O(\psi)$. Our result implies that the nonidentical number of neurons must be $O(1)$ after training. This, in turn, means that the VC dimension of the model is independent of width.